OpenStudy (anonymous):

Assuming a half-life of 1599 years, how many years will be needed for the decay of 15/16 of a given amount of radium-226

OpenStudy (jfraser):

if you need 15/16 to decay, that means only 1/16 will be left. How many times can you cut something in half until you're left with 1/16 of what you started with?

OpenStudy (aaronq):

1 half life: (1/2)^1= 1/2 2 half lives: (1/2)^2= 1/4 3 half lives: (1/2)^3= 1/8